FACTORISATION FREE DECOMPOSITION ALGORITHMS IN DIFFERENTIAL ALGEBRA

EVELYNE HUBERT

Abstract. We present an effective version of Ritt’s algorithm. We apply material of the paper by Boulier, Lazard, Ollivier and Petitot (ISSAC’1995) for which we give new concise proofs. We present original results in constructive algebra that makes the algorithm flexible and simple.